Solve (x+2)(2x-7)=-15
Answer by algebrahouse.com(1659)

Solve (x+2)(2x-7)= -15

(x + 2)(2x - 7) = -15
2x² - 3x - 14 = -15 {used foil method}
2x² - 3x + 1 = 0 {added 15 to both sides}
(2x - 1)(x - 1) = 0 {factored into two binomials}
2x - 1 = 0 or x - 1 = 0 {set each factor equal to 0}
2x = 1 or x = 1 {added 1 to both sides in both equations}
x = 1/2 or x = 1 {divided left equation by 2}
www.algebrahouse.com